An object at a temperature of 400 K. At what wavelength (in nm) does this object radiate the most energy?
Answer #1

Wavelength can be calculated from WEIN'S DISPLACEMENT LAW:-

Which says:-

Wavelength (maxm) * Temperature= 2.893*10^-3

So maxm wavelength= (2.893*10^-3)/400 = 7.2325

{ If you find it helpful please rate positively.
If there's any further queries, or solution is incorrect or presentation is poor, let me know before downvoting. I will reach out to you ASAP.}